IF a cold-water faucet is left on, the sink will fill in 10 mins. If the hot-water faucet is left on, the sink will fill in 12 mins. How long would it take the sink to fill if both faucets are left on?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

filling the facet is 1 job
Cold facuet fills in 10 minutes
So it does 1/10 of the job in 1 minute
Hot facuet fills in 12 minutes
so it does 1/12 of the job in 1 minute
1/10 +1/12 of the job they do together in 1 minute = 22/120
22/120 job in 1 minute
1 job they will do in 120/22 minutes
= 5.45 minutes

Answer by richwmiller(17219)

x/10+x/12=1
12x+10x=120
22x=120
x=120/22
x=60/11
x=5.45 minutes to fill the sink with hot and cold water turned on.